Find all relative extrema. Use the second derivative test where applicable.

(a) f(x)=√x2+1


(b) f(x)=cos x -x


1
Expert's answer
2021-12-03T13:32:51-0500

(a) "f(x)=\sqrt{x^2+1}=(x^2+1)^{1\over 2}"


"f'(x)={1\over 2}(x^2+1)^{-{1\over 2}}.2x={x\over \sqrt {x^2+1}}"

Equate "f'(x)" to to get the value of "x"

"f'(x)={1\over 2}(x^2+1)^{-{1\over 2}}.2x={x\over \sqrt {x^2+1}}=0 \implies x=0"

"f'(x)\lt 0" for "x\lt 0" and "f'(x)\gt0" for "x\gt 0" "\implies" at "x=0" there is a minimum.

Because "f(x)" was decreasing till it reaches a minimum at "x=0" and then started increasing ,because the derivative is positive after "x=0"

(b) "f(x)=cos( x) -x"

"f'(x)=-sin(x)-1"

Equate "f'(x)" to to get the value of "x"

"f'(x)=-sin(x)-1 =0 \implies x={3\pi \over 2}"

"f''(x)=-cos(x)"

"f''({3\pi \over 2})=-cos({3\pi \over 2})=0 \implies" we have an inflection point at "x={3\pi \over 2}"

so,No extrema anywhere